# Bulk plasmons in elemental metals

### v1

@@ Abstract
The spectral properties, momentum dispersion, and broadening of bulk plasmonic excitations of 25 elemental metals are studied from first principles calculations in the random-phase approximation. Spectral band structures are constructed from the resulting momentum- and frequency-dependent inverse dielectric function. We develop an effective analytical representation of the main collective excitations in the dielectric response, by extending our earlier model based on multipole-Padé approximants (MPA) to incorporate both momentum and frequency dependence. With this tool, we identify plasmonic quasiparticle dispersions exhibiting complex features, including non-parabolic energy and intensity dispersions, discontinuities due to anisotropy, and overlapping effects that lead to band crossings and anti-crossings. We also find good agreement between computed results and available experiments in the optical limit. The results for elemental metals establish a reference point that can guide both fundamental studies and practical applications in plasmonics and spectroscopy.
@@ I Introduction
Elemental metals, composed of a single type of metallic atom, are foundational to both modern technology and condensed matter physics [1, 2]. Examples include alkali metals (e.g., Na, K), alkaline earth metals (e.g., Mg, Ca), and transition or noble metals (e.g., Al, Cu, Ag, Au). These materials are characterized by high electrical and thermal conductivity, a strong but ductile metallic bonding, and a conduction band partially filled with delocalized electrons. Delocalized electrons respond collectively to external fields and are responsible for the characteristic optical and electronic properties that distinguish metals from insulators and semiconductors. Thus, understanding the electronic properties of elemental metals, including their collective excitations, is of significant theoretical and practical importance. Metals with a relatively simple electronic band structure, like Al and alkali metals, host delocalized conduction electrons that behave nearly as a free-electron gas [3, 4, 5, 6]. This property, along with a substantial amount of available experimental data, makes them ideal systems for studying fundamental electronic excitations and their underlying many-body interactions.
@@
Collective oscillations of the electron density give rise to plasmonic quasiparticles. Their origin lies in the long-range Coulomb restoring force acting on charge-density fluctuations, which produces well-defined modes in the free-electron gas and, in real materials, dispersions modified by band structure and dielectric screening [7, 8, 9]. Plasmons can be excited by electromagnetic perturbations or incident charged particles, and appear as bulk or surface modes [10, 11]. Plasmons are central in several technological applications, including nanophotonics [12, 13], surface-enhanced spectroscopy [14], energy harvesting [15, 16], and emerging quantum plasmonic technologies [17, 18, 19]. Plasmons can also be engineered by modifying the dielectric environment or the electronic structure, for example through alloying  [20, 21].
@@
Bulk plasmons correspond to longitudinal charge-density oscillations and are therefore optically inactive. They are observed in experiments such as electron energy-loss spectroscopy (EELS) [10, 11], where they appear as well-defined peaks in the measured spectra. Other techniques such as inelastic x-ray scattering and reflection electron energy loss spectroscopy (REELS) are also used to provide complementary insights into plasmonic properties [22, 23, 24, 25, 26, 27, 28]. In simple metals such as Al or Na, the plasmon peak is sharp and symmetric, akin to that of a free-electron gas [5, 10]. In contrast, in noble metals, such as Ag and Au, the occupied $d$-bands introduce inter-band transitions that broaden and shift the plasmon resonance, and give rise to much richer spectra [29, 4]. Such experimental signatures are broadly employed for material characterization, from determining free-carrier concentrations, phases, and thicknesses to exploring nanoplasmonic effects and probing the electronic band structure [27, 11].
@@
Although the dielectric properties of elemental metals in the optical limit have been extensively studied and tabulated [29, 4, 5, 8, 28], properties at finite momentum are not as widely available. In this work, we report the frequency and momentum dependent inverse dielectric functions of 25 elemental metals computed from first principles within the random phase approximation (RPA) level of the many-body theory, using Kohn–Sham Bloch (KS) states obtained from density functional theory (DFT). We also provide an effective analytical representation of the main collective excitations in the inverse dielectric function, by generalizing our previous model based on multipole-Padé approximants (MPA) [30, 31, 32] through the inclusion of both momentum and frequency dependence. The generalized MPA model in two variables, MPA($\mathbf{q}$}), is used to describe the main spectral properties of such excitations.
@@ II.1 Theoretical framework
The complex dielectric function $\varepsilon$ governs the response of a material to external electromagnetic fields. It depends on momentum $\mathbf{q}$} and frequency $\omega$. In the optical or long-wavelength limit, ${\mathbf{q}}\to 0$, the dielectric function of the free-electron gas reduces to the Drude form, characterized by a single pole [33, 4, 34, 5, 35]. Within the RPA approximation, Lindhard theory extends this classical Drude model to include the $\mathbf{q}$} dependence [9, 8], providing a first-order quantum-mechanical description of the free-electron response. The dielectric response in real materials, often deviates from the homogeneous fee-electron behavior, and is therefore much richer. The RPA dielectric function can be computed by first-principle calculations using KS-DFT states. In such an approach, the independent-particle microscopic polarizability is given by
@@
In metallic systems, Eq. (1) encompasses intra- and inter-band transitions that mix in the Dyson equation of Eq. (2). We can thus separate both contributions to the interacting polarizability as
$\chi^{\mathrm{intra}}$ $=\chi-\chi^{\mathrm{inter}}$ (4)
@@
where $Y_{\mathrm{intra}}\equiv v\chi^{\mathrm{intra}}$ and $Y_{\mathrm{inter}}\equiv v\chi^{\mathrm{inter}}$. Even if $\chi^{\mathrm{intra}}$ vanishes in the ${\mathbf{q}}\to 0$ limit, $Y_{\mathrm{intra}}$ remains finite due to the divergence of the Coulomb potential $v$. Therefore, the plasmonic quasiparticles observed in $Y({\mathbf{q}},\omega)$ at any value of $\mathbf{q}$} are a mixture of intra- and inter-band contributions.
An important property of the imaginary part of $Y$, or loss function, is the f-sum rule, which can be used to define the intra-band, inter-band, and plasma frequencies:
$\omega_{\mathrm{intra}}^{2}({\mathbf{q}})$ $\equiv\frac{2}{\pi}\int_{0}^{\infty}d\omega\,\omega|\mathrm{Im}[Y_{\mathrm{intra}}({\mathbf{q}},\omega)]|$ (7)
$\omega_{\mathrm{inter}}^{2}({\mathbf{q}})$ $\equiv\frac{2}{\pi}\int_{0}^{\infty}d\omega\,\omega|\mathrm{Im}[Y_{\mathrm{inter}}({\mathbf{q}},\omega)]|$ (8)
@@
The interacting $\chi$, and hence also $Y$, has a pole structure similar to that of $\chi_{0}$ in Eq. (1). However, the large number of single-particle transitions $\Omega^{\mathrm{KS}}_{vc\mathbf{k}\mathbf{q}}$ are mixed in the Dyson equation in Eq. (2), typically resulting in much simpler envelope functions [31]. For many systems, $Y$ in the optical limit ${\mathbf{q}}\to 0$ resembles the response of the free-electron gas [36, 37, 8], and can therefore be approximated by a single plasmon-pole model (PPA):
@@
where the pole is located at the plasmon energy, $\mathrm{Re}[\Omega_{0}]=\omega_{\mathrm{pl}}$. In the free-electron gas the plasmon energy is related to the electronic density, $\rho_{e}$, as (Gaussian units)
@@
$Y^{\mathrm{PPA}}({\mathbf{q}},\omega)=\frac{2R({\mathbf{q}})\Omega({\mathbf{q}})}{\omega^{2}-[\Omega({\mathbf{q}})]^{2}}\,.$ (12)
where the pole has a quadratic dispersion according to the Drude model, $\Omega(q)\approx\Omega_{0}(1+\Omega^{\prime\prime}q^{2}/2)$, with an effective electron mass factor $\Omega^{\prime\prime}$. The dispersion of the spectral weight is linked to the position of the pole by the f-sum rule of Eq. (9):
@@ II.2 Simple analytical representation: the multipole approach (MPA)
The RPA approximation is at the core of ab initio many-body theories beyond DFT, such as $GW$; however, expressing the operators in terms of single-particle excitations, as in the Lehmann representation, is both computationally impractical and makes difficult the physical interpretation of collective features. Alternatively, analytical models simplifying the frequency dependence of the dielectric function, starting from PPA, have been used to reduce the computational cost of first-principle calculations. However, the use of PPA is subject to accuracy limitations [30, 31]. More complex models, such as Padé approximants and many-pole schemes, have often been used to fit and interpret individual experimental spectra [46, 47, 48, 49, 50, 25, 51]. Within ab initio $GW$ calculations, effective MPA representations have been shown to provide an efficient description of the dielectric response, achieving an accuracy comparable to that of full-frequency approaches [30, 31, 32]. Such models generalize the PPA inverse dielectric response of Eq. (12), as follows
@@
$\begin{split}R_{p}(q)&\approx R_{p}(1+R^{\prime}_{p}q+R^{\prime\prime}_{p}q^{2}/2+R^{\prime\prime\prime}_{p}q^{3}/6)\\ \Omega_{p}(q)&\approx\Omega_{p}(1+\Omega^{\prime}_{p}q+\Omega^{\prime\prime}_{p}q^{2}/2+\Omega^{\prime\prime\prime}_{p}q^{3}/6),\end{split}$ (15)
where $R_{p}$ and $\Omega_{p}$ are the residues and pole positions in the optical limit $q\to 0$ along the selected direction, and the respective pairs $R^{\prime}_{p}$ and $\Omega^{\prime}_{p}$, $R^{\prime\prime}_{p}$ and $\Omega^{\prime\prime}_{p}$, and $R^{\prime\prime\prime}_{p}$ and $\Omega^{\prime\prime\prime}_{p}$ are the corresponding linear, quadratic, and cubic coefficients of the polynomial expansions. Although other $q$ dependencies can be explored, a third-order polynomial is shown to be sufficiently flexible to obtain accurate representations with relatively few poles, from 2 to 15, for all the systems studied here. Therefore, the generalized MPA($\mathbf{q}$}) model results in a quite compact representation of the momentum and frequency dependent $Y({\mathbf{q}},\omega)$, with the total number of parameters given by $8\times n_{Y}$. All the parameters are determined in a non-linear fit with physical constraints on the position of the poles (see Refs. [30, 31]):
@@
However, this is strictly respected only when the integration interval over energy in Eq. (9) is sufficiently large. Therefore, we do not explicitly impose it for small energy ranges.
@@ II.3 Reference data for the set of elemental metals
The dielectric properties of metals in the optical limit are widely studied and reported, e.g., in several handbooks and websites [54, 4]. However, the frequency dependent dielectric function is not always available and different sources can be inconsistent with respect to precision, energy range, etc. Experimental techniques such as EELS only measure the imaginary part of the response and need careful post-processing to retrieve the inverse dielectric function, which is nontrivial at finite momentum [55].
@@
A complete list of the 25 elemental metals studied here is shown in Table 1. The table includes the atomic number, the electronic configuration of valence electrons, the symmetry group, the experimental lattice parameters [70], and the volume of the unit cell. The computed intra- and inter-band frequencies are also provided.
@@ II.4 Computational details
The crystalline structure of elemental metals is well characterized experimentally and theoretically [71]. In this work, we use the experimental geometries and lattice parameters listed in Table 1. DFT calculations were performed using the plane wave implementation of the Quantum Espresso package [72, 73] with the Perdew-Burke-Ernzerhof (PBE) variant of the GGA functional [74]. We adopted the norm-conserving optimized Vanderbilt pseudopotentials of Ref. [75], with a kinetic plane wave energy cutoff of 70 Ry. In the DFT calculations, the Brillouin zone was sampled with a $36\times 36\times 36$ Monkhorst-Pack grid for cubic, and $36\times 36\times 24$ for hexagonal materials. The calculations of dielectric spectra within RPA were performed with the yambo code [76, 77].

### v2

@@ Abstract
The spectral properties, momentum dispersion, and broadening of bulk plasmonic excitations of 26 elemental metals are studied from first principles calculations in the random-phase approximation. Spectral band structures are constructed from the resulting momentum- and frequency-dependent inverse dielectric function. We develop an effective analytical representation of the main collective excitations in the dielectric response, extending our earlier model based on multipole-Padé approximants (MPAs) to incorporate both momentum and frequency dependence [MPA($\mathbf{q}$})]. With this representation, we identify plasmonic quasiparticle dispersions exhibiting complex features, including non-parabolic energy and intensity dispersions, discontinuities due to anisotropy, and overlapping effects that lead to band crossings and anti-crossings. Comparing with available experimental data, mainly in the optical limit, we find good agreement with the computed spectra. The results for elemental metals and their effective MPA($\mathbf{q}$}) representation establish a reference point that can guide both fundamental studies and practical applications in plasmonics and spectroscopy.
@@ I Introduction
Elemental metals, composed of a single type of metallic atom, are foundational to both modern technology and condensed matter physics [7, 39]. Examples encompass alkali metals (e.g., Na, K), alkaline earth metals (e.g., Mg, Ca), and noble (e.g., Al, Cu, Ag, Au) or other transition metals (e.g., Cr, Ti, Fe, Ni). These materials are characterized by high electrical and thermal conductivity, a strong but ductile metallic bonding, and a conduction band partially filled with delocalized electrons. Delocalized electrons respond collectively to external fields and are responsible for the characteristic optical and electronic properties that distinguish metals from insulators and semiconductors. Thus, understanding the electronic properties of elemental metals, including their collective excitations, is of significant theoretical and practical importance. Metals with a relatively simple electronic band structure, such as Al and alkali metals, host delocalized conduction electrons that behave nearly as a free-electron gas [54, 60, 51, 49]. This property, along with a substantial amount of available experimental data, makes them ideal systems for studying fundamental electronic excitations and their underlying many-body interactions.
@@
Collective oscillations of the electron density give rise to plasmonic quasiparticles, or plasmons. Their origin lies in the long-range Coulomb restoring force acting on charge-density fluctuations, which produces well-defined modes in the free-electron gas, while in real materials, the plasmonic dispersions are modulated by their underlying electronic band structure and dielectric screening [63, 24, 50]. Plasmons can be excited by electromagnetic perturbations or incident charged particles, and appear as bulk or surface modes [66, 18]. Plasmons are central in several technological applications, including nanophotonics [11, 36], surface-enhanced spectroscopy [57], energy harvesting [8, 17], and emerging quantum plasmonic technologies [32, 25, 10]. Plasmonic properties can also be engineered by modifying the dielectric environment or the electronic structure, for example through alloying  [56, 59].
@@
Bulk plasmons correspond to longitudinal charge-density oscillations and are therefore optically inactive. They can be probed experimentally by methods such as electron energy-loss spectroscopy (EELS) [66, 18], where they appear as well-defined peaks in the measured spectra. Other techniques such as inelastic x-ray scattering and reflection electron energy loss spectroscopy (REELS) can also provide complementary insights into plasmonic properties [71, 70, 69, 72, 84, 21, 9]. In simple metals such as Al or Na, the plasmon peak is sharp and symmetric, akin to that of a free-electron gas [51, 66]. In contrast, in noble metals, such as Ag and Au, the occupied $d$-bands introduce inter-band transitions that broaden and shift the plasmon resonance, and give rise to much richer spectra [34, 60]. Such experimental signatures are broadly employed for material characterization, from determining free-carrier concentrations, crystalline phases, and thicknesses to exploring nanoplasmonic effects and probing the electronic band structure [21, 18, 46].
@@
Although the dielectric properties of elemental metals in the optical limit have been extensively studied and tabulated [34, 60, 51, 24, 9, 64], properties at finite momentum are not as widely available. For gapless metallic systems, strong electronic screening suppresses bound electron–hole (excitonic) effects, making the random-phase approximation (RPA) a well-established and widely used framework for describing bulk plasmon excitations [24, 49]. In this work, we report the frequency and momentum dependent inverse dielectric functions of 26 elemental metals computed from first principles within the RPA level of the many-body theory, using Kohn–Sham Bloch (KS) states obtained from density functional theory (DFT). We also provide an effective analytical representation of the main collective excitations in the inverse dielectric function, by generalizing our previous model based on multipole-Padé approximants (MPA) [45, 47, 44] to include momentum alongside frequency dependence. The generalized MPA model in two variables, MPA($\mathbf{q}$}), is used to describe the main spectral properties of such excitations. It also shows potential to develop efficient methods to further reduce the computational cost of first-principles calculations such as $GW$ and BSE [49, 46], with respect to the original MPA method [45].
@@ II.1 Theoretical framework
The complex dielectric function $\varepsilon$ governs the response of a material to external electromagnetic fields. It depends on momentum $\mathbf{q}$} and frequency $\omega$. In the optical, or long-wavelength limit, ${\mathbf{q}}\to 0$, the dielectric function of the free-electron gas reduces to the Drude form, characterized by a single plasmon pole [6, 60, 75, 51, 42]. Within the RPA approximation, Lindhard theory extends the classical Drude model to include the $\mathbf{q}$} dependence [50, 24], providing a first-order quantum-mechanical description of the free-electron response. The dielectric response of real materials, often deviates from that of the homogeneous free-electron behavior, and is often much richer. The RPA dielectric function can be computed from first principles using KS-DFT states to evaluate the independent-particle microscopic polarizability, as follows
@@
In metallic systems, $\chi_{0}$ encompasses both intra- and inter-band transitions, which in Eq. (1) correspond to $n=m$ and $n\neq m$ respectively, and thus $\chi_{0}=\chi_{0}^{\mathrm{intra}}+\chi_{0}^{\mathrm{inter}}$. In the interacting polarizability $\chi$, such contributions mix according to the Dyson equation of Eq. (2) and cannot be simply splitted. In order to define intra- and inter-band frequencies from the interacting $\chi$, we can then take the case of semiconductors, that have only inter-band contributions, to define the following terms
$\chi^{\mathrm{intra}}$ $=\chi-\chi^{\mathrm{inter}},$ (4)
@@
where $Y_{\mathrm{intra}}\equiv v\chi^{\mathrm{intra}}$ and $Y_{\mathrm{inter}}\equiv v\chi^{\mathrm{inter}}$. Since $\chi_{0}^{\mathrm{intra}}$ vanishes in the ${\mathbf{q}}\to 0$ limit, so does $\chi^{\mathrm{intra}}$. However, $Y_{\mathrm{intra}}$ remains finite due to the divergence of the Coulomb potential $v$. Therefore, in metals, the plasmonic quasiparticles observed in $Y({\mathbf{q}},\omega)$ at any value of $\mathbf{q}$} emerge from a mixture of intra- and inter-band contributions.
An important property of the imaginary part of $Y$, or the loss function $\mathrm{Im}[-\varepsilon^{-1}]$, is the f-sum rule, which is conserved in the inversion of the Dyson equation (see Supplemental materials of Ref. [47]) and thus can be used to define corresponding intra-band, inter-band, and plasma frequencies:
$\omega_{\mathrm{intra}}^{2}({\mathbf{q}})$ $\equiv\frac{2}{\pi}\int_{0}^{\infty}d\omega\,\omega|\mathrm{Im}[Y_{\mathrm{intra}}({\mathbf{q}},\omega)]|,$ (7)
$\omega_{\mathrm{inter}}^{2}({\mathbf{q}})$ $\equiv\frac{2}{\pi}\int_{0}^{\infty}d\omega\,\omega|\mathrm{Im}[Y_{\mathrm{inter}}({\mathbf{q}},\omega)]|,$ (8)
@@
The interacting $\chi$, and hence also $Y$, has a pole structure similar to that of $\chi_{0}$ in Eq. (1). However, the large number of single-particle transitions $\Omega^{\mathrm{KS}}_{nm\mathbf{k}\mathbf{q}}$ are mixed in the Dyson equation in Eq. (2), giving rise to plasmonic quasiparticles and thus resulting in much simpler envelope functions [47]. For many systems, $Y$ in the optical limit ${\mathbf{q}}\to 0$ resembles the response of the free-electron gas [29, 20, 24], and can therefore be approximated by a single plasmon-pole model (PPA):
@@
where the pole is located at the plasmon energy, $\mathrm{Re}[\Omega_{0}]=\omega_{\mathrm{pl}}$. In the free-electron gas the plasmon energy is related to the electronic density, $\rho_{e}$, as (in Gaussian units)
@@
$Y^{\mathrm{PPA}}({\mathbf{q}},\omega)=\frac{2R({\mathbf{q}})\Omega({\mathbf{q}})}{\omega^{2}-[\Omega({\mathbf{q}})]^{2}},$ (12)
where the pole has a quadratic dispersion according to the Drude/Lindhard model, $\Omega(q)\approx\Omega_{0}(1+\Omega^{\prime\prime}q^{2}/2)$, with an effective electron mass factor $\Omega^{\prime\prime}$. The dispersion of the spectral weight is linked to the position of the pole by the f-sum rule of Eq. (9):
@@ II.2 Simple analytical representation: The multipole approach (MPA)
The RPA approximation is at the core of ab initio many-body theories beyond DFT, such as $GW$; however, expressing the operators in terms of single-particle excitations, as in the Lehmann representation of Eq. (1), is both computationally impractical and hinders the physical interpretation of collective features. Within first principles approaches, analytical models simplifying the frequency dependence of the dielectric function, starting from the (single-pole) PPA, have been used to reduce the computational cost. However, the use of PPA is subject to accuracy limitations [45, 47]. More complex models, such as Padé approximants and many-pole schemes, have often been used to fit and interpret individual experimental spectra [43, 33, 76, 37, 38, 72, 14]. Within ab initio $GW$ calculations, effective MPA representations have been shown to provide an efficient description of the dielectric response, achieving an accuracy comparable to that of full-frequency approaches [45, 47, 44]. Such models generalize the PPA inverse dielectric response of Eq. (12) as
@@
$\begin{split}R_{p}(q)&\approx R_{p}(1+R^{\prime}_{p}q+R^{\prime\prime}_{p}q^{2}/2+R^{\prime\prime\prime}_{p}q^{3}/6),\\ \Omega_{p}(q)&\approx\Omega_{p}(1+\Omega^{\prime}_{p}q+\Omega^{\prime\prime}_{p}q^{2}/2+\Omega^{\prime\prime\prime}_{p}q^{3}/6),\end{split}$ (15)
where $R_{p}$ and $\Omega_{p}$ are the residues and pole positions in the optical limit $q\to 0$ along the selected direction, and the respective pairs $R^{\prime}_{p}$ and $\Omega^{\prime}_{p}$, $R^{\prime\prime}_{p}$ and $\Omega^{\prime\prime}_{p}$, and $R^{\prime\prime\prime}_{p}$ and $\Omega^{\prime\prime\prime}_{p}$ are the corresponding linear, quadratic, and cubic coefficients of the polynomial expansions. Although other $q$ dependencies can be explored, we find a third-order polynomial to be sufficiently flexible to obtain accurate representations with relatively few poles, between 2 and 16, for all the systems studied here. Therefore, the generalized MPA($\mathbf{q}$}) model provides a fairly compact representation of the momentum and frequency dependent $Y({\mathbf{q}},\omega)$, with the total number of parameters given by $8\times n_{Y}$. All the parameters are determined in a non-linear fit with physical constraints on the position of the poles, satisfying the time-ordering in a finite energy interval (see Refs. [45, 47]):
@@
However, this is strictly respected only when the integration interval over energy in Eq. (9) is sufficiently large. Therefore, we do not explicitly impose it in any of the present calculations, since they are done for limited energy ranges.
@@ II.3 Reference data for the set of elemental metals
The dielectric properties of metals in the optical limit are widely studied and reported, e.g., in several handbooks and websites [1, 60, 64]. However, the frequency dependent dielectric function is not always available and different sources can be inconsistent with respect to precision, energy range, etc. Experimental techniques such as EELS only measure the imaginary part of the response and need careful post-processing to retrieve the inverse dielectric function, which is nontrivial at finite momentum [46].
@@
A complete list of the 26 elemental metals studied here is shown in Table 1. The table includes the atomic number, the electronic configuration of valence electrons, the symmetry group, the experimental lattice parameters [2], and the volume of the unit cell. The computed intra- and inter-band frequencies are also provided.
@@ II.4 Computational details
The crystalline structure of elemental metals is well characterized experimentally and theoretically [81]. In this work, we use the experimental geometries and lattice parameters listed in Table 1. DFT calculations were performed using the Quantum Espresso [23, 22] plane wave package with the Perdew-Burke-Ernzerhof variant of the GGA functional [61]. We adopted the norm-conserving optimized Vanderbilt pseudopotentials of Ref. [28], with a kinetic plane wave energy cutoff of 70 Ry. In the DFT calculations, the Brillouin zone was sampled with a $36\times 36\times 36$ Monkhorst-Pack grid for cubic, and $36\times 36\times 24$ for hexagonal materials. Metallic occupations were treated using a smearing technique, employing ordinary Gaussian spreading with a smearing width of 0.01 Ry. Spin polarization was included in the DFT calculations of ferromagnetic elemental metals (Fe, Co, Ni). The calculations of dielectric spectra within RPA were performed with the yambo code [52, 68]. Eq. (1) is evaluated with a damping parameter of $\delta=0.1$ eV and a kinetic plane wave energy cutoff of 5 Ry.